# Hallucination Detection in Foundation Models for Decision-Making: A Flexible Definition and Review of the State of the Art

### v1

@@ Abstract
Autonomous systems are soon to be ubiquitous, from manufacturing autonomy to agricultural field robots, and from health care assistants to the entertainment industry. The majority of these systems are developed with modular sub-components for decision-making, planning, and control that may be hand-engineered or learning-based. While these existing approaches have been shown to perform well under the situations they were specifically designed for, they can perform especially poorly in rare, out-of-distribution scenarios that will undoubtedly arise at test-time. The rise of foundation models trained on multiple tasks with impressively large datasets from a variety of fields has led researchers to believe that these models may provide “common sense” reasoning that existing planners are missing. Researchers posit that this common sense reasoning will bridge the gap between algorithm development and deployment to out-of-distribution tasks, like how humans adapt to unexpected scenarios. Large language models have already penetrated the robotics and autonomous systems domains as researchers are scrambling to showcase their potential use cases in deployment. While this application direction is very promising empirically, foundation models are known to hallucinate and generate decisions that may sound reasonable, but are in fact poor. We argue there is a need to step back and simultaneously design systems that can quantify the certainty of a model’s decision, and detect when it may be hallucinating. In this work, we discuss the current use cases of foundation models for decision-making tasks, provide a general definition for hallucinations with examples, discuss existing approaches to hallucination detection and mitigation with a focus on decision problems, and explore areas for further research in this exciting field.
1 Introduction
@@
It is an exciting time to be a machine learning and robotics researcher. A great deal of progress has been made in the last decade and a half with regards to the efficacy and efficiency of models for perception, decision-making, planning, and control Soori et al. (2023); Janai et al. (2020). Broadly speaking, approaches to these problems fall under one of two umbrellas: hand-engineered model-based systems and data-driven learning-based models Formentin et al. (2013). With some deployment scenario in mind, developers may hand-engineer rules Hayes-Roth (1985) or tune a controller Borase et al. (2021) to be tested, or in the case of learning-based models, collect training data and craft some reward function to fit a model to an objective, given said data Henderson et al. (2018). In practice, these methods work particularly well in the scenarios that they were specifically designed and trained for, but may produce undesirable results in previously unseen out-of-distribution cases Wen et al. (2023). Designers may choose to add more rules, re-tune their controller, fine-tune their model to a more representative dataset, fix the reward function to handle edge cases, or even add a detector (which may itself be rule-based or data-driven) at test-time to identify out-of-distribution scenarios before calling on the decision-maker Singer and Cohen (2021); Schreiber et al. (2023); Chakraborty et al. (2023). However, even with these changes, there will always be other situations that designers had not previously considered which will come about during deployment, leading to sub-optimal performance or critical failures. Furthermore, the modifications made to the model may have unforeseen effects at test-time like undesired conflicting rules Ekenberg (2000) or catastrophic forgetting of earlier learned skills Kemker et al. (2018).
@@
Informally, classical methods and data-driven approaches lack some form of common sense that humans use to adapt in unfamiliar circumstances Fu et al. (2023a). More recently, researchers are exploring the use of large (visual) language models, L(V)LMs, to fill this knowledge gap Cui et al. (2024). These models are developed by collecting and cleaning an enormous natural language dataset, pre-training to reconstruct sentences on said dataset, fine-tuning on specific tasks (e.g., question-answering), and applying human-in-the-loop reinforcement learning to produce more reasonable responses Achiam et al. (2023). Even though these models are another form of data-driven learning that attempt to maximize the likelihood of generated text conditioned on a given context, researchers have shown that they have the ability to generalize to tasks they have not been trained on, and reason about their decisions. As such, these foundation models are being tested in tasks like simulated decision-making Huang et al. (2024b) and real-world robotics Zeng et al. (2023) to take the place of perception, planning, and control modules. Even so, foundation models are not without their limitations. Specifically, these models have a tendency to hallucinatei.e., generate decisions or reasoning that sound plausible, but are in fact inaccurate or would result in undesired effects in the world. This phenomenon has led to the beginning of a new research direction that attempts to detect when L(V)LMs hallucinate so as to produce more trustworthy and reliable systems. Before these large black-box systems are applied in safety-critical situations, there need to be methods to detect and mitigate hallucinations. Thus, this survey collects and discusses current hallucination mitigation techniques for foundation models in decision-making tasks, and presents potential research directions.
@@
Existing surveys particularly focus on presenting methods for hallucination detection and mitigation in question-answering (QA) Ji et al. (2023); Rawte et al. (2023); Zhang et al. (2023d); Ye et al. (2023) or object detection tasks Li et al. (2023c). There are also other works that provide examples of current use cases of L(V)LMs in autonomous vehicles Yang et al. (2023b) and robotics Zeng et al. (2023); Zhang et al. (2023a). Wang et al. (2023a) perform a deep analysis of the trustworthiness of a variety of foundation models and Chen and Shu (2024) provide a taxonomy of hallucinations within LLMs, but both exclude applications to general decision problems. To the best of our knowledge, we are the first to propose a general definition of hallucinations that can be flexibly tuned to any particular deployment setting, including commonly found applications to QA or information retrieval, and more recent developments in planning or control. Furthermore, there is no existing work that summarizes state of the art methods for hallucination detection and mitigation approaches within decision-making and planning tasks.
@@
In the remainder of this work, we discuss the current uses of foundation models for decision-making tasks in Section 2, define and provide examples of hallucinations in Section 3, identify current detection methods and where they are evaluated in Sections 4 and 5 respectively, and explore possible research directions in Section 6.
2 Foundation Models Making Decisions
@@
Originally coined by Bommasani et al. (2022), the term foundation models refers to models that are “trained on broad data at scale such that they can be adapted to a wide range of downstream tasks.” This approach is in contrast to works that design and train models on a smaller subset of data for the purpose of being deployed to a specific task Yang et al. (2024). The key difference is that foundation models undergo a pre-training procedure on a large-scale dataset containing information from a variety of possible deployment fields, through which they are expected to learn more general features and correspondences that may be useful at test-time on a broader set of tasks Zhou et al. (2023); Zhao et al. (2023). Examples of existing pre-trained foundation models span language Devlin et al. (2019); Brown et al. (2020); Touvron et al. (2023a), vision Caron et al. (2021); Oquab et al. (2024); Kirillov et al. (2023), and multi-modal Radford et al. (2021); Achiam et al. (2023) inputs. In this section, we give a brief overview of existing use cases for foundation models in robotics, autonomous vehicles, and other decision-making systems. We also succinctly point out hallucinations found in these works and leave a lengthier discussion in Section 3.2. Readers should refer to works from Yang et al. (2023b)Zeng et al. (2023), and Zhang et al. (2023a) for a deeper review of application areas.
2.1 Autonomous Driving
@@
For the autonomous vehicle domain, researchers have formulated the use of language foundation models as a fine-tuning and prompt engineering problem. An external sub-system is usually designed with (1) a perception module to process signals from raw sensors, (2) a memory bank of prior important experiences and its corresponding similarity function to find alike scenarios, and (3) a prompt generator to convert current sensor data and relevant memories into natural language that can be input to the foundation model. Currently, works either fine-tune LLMs with a few examples, or directly apply the model in a zero-shot manner, on a QA task with driving related questions. By framing the task in a QA form, researchers have been able to provide context to the L(V)LM to probe for high-level natural language decisions Wen et al. (2023, 2024), path planning Mao et al. (2023); Sima et al. (2023), vehicle tracking and trajectory prediction Wu et al. (2023b); Keysan et al. (2023), descriptions of the surroundings of the vehicle Chen et al. (2023c); Xu et al. (2024), and low-level control Liu et al. (2023).
@@
Wen et al. (2024) propose DiLu, a framework consisting of reasoning, reflection, and memory modules that support an LLM in producing high-level decisions for autonomous driving, and they test their method within a driving simulator environment. Specifically, the reasoning module views the current observation of the vehicle, queries the memory module for any similar situations that were encountered in the past, and converts the experience into a prompt, which is input to the LLM. The prompt is formatted such that it elicits chain-of-thought reasoning Wei et al. (2022) from the LLM, which is shown to improve the accuracy of the model. The generated text output by the LLM is summarized by the reflection module, and is used to update the memory bank of experiences. A separate decision decoder model converts the summary into a discrete high-level decision (e.g., idle, turn right, accelerate, etc.) to take in the simulator. The same authors have also experimented with prompting the LVLM model, GPT-4V, in a zero-shot manner to describe the surroundings of the vehicle, take high-level decisions, and explain why it believes it would be a good action to take Wen et al. (2023). They find that GPT-4V is capable of identifying safety-critical scenarios and suggests driving more conservatively in those situations. However, like other existing vision models Suganuma and Yoneda (2022), it has difficulty in detecting traffic light states. We discuss other hallucination examples in traffic scenarios in Section 3.2.
@@
Agent-Driver from Mao et al. (2023) utilizes a tool library of functions that communicate with neural modules that are responsible for object detection, trajectory prediction, occupancy estimation, and mapping. The LLM is asked to reason about what information would be helpful to plan a path of the ego vehicle, and calls on functions from the tool library to iteratively build up relevant context. Like Wen et al. (2024), the authors use a memory bank of prior driving experiences to bolster the context provided to the LLM. With this context, the LLM predicts a suitable path for the ego vehicle to follow. If a collision is detected between the predicted trajectory and surrounding objects in the scene, the LLM undergoes self-reflection, like Reflexion Shinn et al. (2023), another hallucination mitigation technique, to fine-tune its prediction. Through a short study, the authors test the frequency of invalid, hallucinated outputs from the model, and find that their self-reflection approach results in zero invalid generations at test-time. Sima et al. (2023) build up context before predicting a path for the ego vehicle by asking a VLM questions about its perception of surrounding vehicles, predicting their behaviors, planning high-level ego vehicle decisions, converting to lower level discrete actions, and finally, estimating a coordinate-level trajectory. Their method, DriveLM-Agent, predicts paths from raw images in an end-to-end manner using a multi-modal approach, whereas Agent-Driver requires sensor modules to process context separately.
@@
Works using foundation models for generating scene descriptions given multi-modal information frame the task as a QA problem. For example, Chen et al. (2023c) use a reinforcement learning (RL) agent pre-trained on the driving task in simulation to collect a dataset containing the vehicle’s state, environment observation (assumed to be ground truth from simulator,) low-level action, and the ego’s percentage of attention placed on different surrounding agents (if using an attention-based network). The authors ask ChatGPT OpenAI (2022) to act as a professional driving instructor who generates questions and answers given general driving rules and the vectorized description of the scene from the RL agent. Then, combining a pre-trained LLM with a vector embedder and former model, the architecture is trained end-to-end to answer questions from the scene. Examples of questions the architecture is posed at test-time include, “What objects are you observing,” “How are you going to drive in this situation and why,” and, “What are the best tourist spots in London” (the last of which is considered out of scope by the driving model). The authors acknowledge that the LLM may generate undesired hallucinated outputs during deployment, so they augment their training instruction dataset with out of scope questions that the model should learn to refuse to answer. DriveGPT4, proposed by Xu et al. (2024), is a multi-modal LLM using LLaMA 2 Touvron et al. (2023b) which encodes video sequences with an encoder model, and projects question and video embeddings into a text form to be input to the LLM. A decoder model converts the tokenized output of the LLM into a low-level action with a corresponding high-level reason to take the action. Like Chen et al. (2023c), the authors collect a driving QA dataset to align and fine-tune the architecture to driving tasks.
@@
Liu et al. (2023) tackle low-level control within unsignalized intersections by formulating the problem as a multi-task decision-making process. Specifically, they train expert policies with RL to perform individual control tasks (e.g., turning left, going forward, and turning right), with which they collect an expert demonstration dataset. An LLM model, GPT-2 Radford et al. (2019), is fine-tuned to predict next actions given past trajectories of state and action pairs from the expert dataset. The authors showcase that their method, MTD-GPT, is able to achieve higher success rates across each of the control tasks over the original RL policies alone, showcasing the promise of foundation models to leverage their general knowledge in a multi-task problem setting.
2.2 Robotics

### v2

@@ Abstract.
Autonomous systems are soon to be ubiquitous, spanning manufacturing, agriculture, healthcare, entertainment, and other industries. Most of these systems are developed with modular sub-components for decision-making, planning, and control that may be hand-engineered or learning-based. While these approaches perform well under the situations they were specifically designed for, they can perform especially poorly in out-of-distribution scenarios that will undoubtedly arise at test-time. The rise of foundation models trained on multiple tasks with impressively large datasets has led researchers to believe that these models may provide “common sense” reasoning that existing planners are missing, bridging the gap between algorithm development and deployment. While researchers have shown promising results in deploying foundation models to decision-making tasks, these models are known to hallucinate and generate decisions that may sound reasonable, but are in fact poor. We argue there is a need to step back and simultaneously design systems that can quantify the certainty of a model’s decision, and detect when it may be hallucinating. In this work, we discuss the current use cases of foundation models for decision-making tasks, provide a general definition for hallucinations with examples, discuss existing approaches to hallucination detection and mitigation with a focus on decision problems, present guidelines, and explore areas for further research in this exciting field.
Foundation Models, Decision-Making, Hallucination Detection and Mitigation, Survey
This work is supported by the Sponsor Office of Naval Research under Grant No.: Grant #N00014-23-1-2651.
copyright: ccjournal: CSURjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3716846ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computing methodologies Natural language generationccs: Computing methodologies Machine learning approaches
1. Introduction
@@
A great deal of progress has been made in the last decade and a half with regards to the efficacy and efficiency of models for perception, decision-making, planning, and control (Soori et al., 2023; Janai et al., 2020). Broadly speaking, approaches to these problems fall under one of two umbrellas: hand-engineered model-based systems and data-driven learning-based models (Formentin et al., 2013). With some deployment scenario in mind, developers may hand-engineer rules (Hayes-Roth, 1985) or tune a controller (Borase et al., 2021) to be tested, or in the case of learning-based models, collect training data and craft some reward function to fit a model to an objective, given said data (Henderson et al., 2018). In practice, these methods work particularly well in the scenarios that they were specifically designed and trained for, but may produce undesirable results in previously unseen out-of-distribution cases (Wen et al., 2024b). Designers may choose to add more rules, re-tune their controller, fine-tune their model to a more representative dataset, fix the reward function to handle edge cases, or even add a detector (which may itself be rule-based or data-driven) at test-time to identify out-of-distribution scenarios before calling on the decision-maker (Singer and Cohen, 2021; Schreiber et al., 2023; Chakraborty et al., 2023). However, even with these changes, there will always be other situations that designers had not previously considered which will come about during deployment, leading to sub-optimal performance or critical failures. Furthermore, the modifications made to the model may have unforeseen effects at test-time like undesired conflicting rules (Ekenberg, 2000) or catastrophic forgetting of earlier learned skills (Kemker et al., 2018).
@@
Informally, classical methods and data-driven approaches lack some form of common sense that humans use to adapt in unfamiliar circumstances (Fu et al., 2024a). More recently, there has been work towards developing multi-modal large language models, which take inputs in the form of images, videos, audio, and text, to tackle more complex language understanding and reasoning tasks (Bommasani et al., 2022; Bai et al., 2024). These models are developed by collecting and cleaning an enormous natural language dataset, pre-training to reconstruct sentences on said dataset, fine-tuning on specific tasks (e.g., question-answering), and applying human-in-the-loop reinforcement learning to produce more reasonable responses (Achiam et al., 2023). Even though these models are another form of data-driven learning that attempt to maximize the likelihood of generated text conditioned on a given context, researchers have shown that they have the ability to generalize to tasks they have not been trained on, and reason about their decisions. Many researchers are specifically exploring the use of large (visual) language models, L(V)LMs, to fill the knowledge gap found in earlier works (Cui et al., 2024). As such, these foundation models are being tested in tasks like simulated decision-making (Huang et al., 2024b) and real-world robotics (Zeng et al., 2023) to take the place of perception, planning, and control modules. Even so, foundation models are not without their limitations. Specifically, these models have a tendency to hallucinatei.e., generate decisions or reasoning that sound plausible, but are in fact inaccurate or would result in undesired effects in the world (Ji et al., 2023; Rawte et al., 2023; Dai et al., 2023). This phenomenon has led to the beginning of a new research direction that attempts to detect when L(V)LMs hallucinate so as to produce more trustworthy and reliable systems. Before these large black-box systems are applied in safety-critical situations, there need to be methods to detect and mitigate hallucinations. Thus, this survey collects and discusses current hallucination mitigation techniques for foundation models in decision-making tasks, and presents potential research directions.
@@
Existing surveys particularly focus on presenting methods for hallucination detection and mitigation in question-answering (QA) (Ji et al., 2023; Rawte et al., 2023; Zhang et al., 2023b; Ye et al., 2023) or object detection tasks (Li et al., 2023b). For example, Ji et al. (2023) summarize metrics and hallucination detection and mitigation methods for abstractive summarization, dialogue generation, generative question-answering, data-to-text generation, and machine translation for natural language generation models in general. Ye et al. (2023) and Rawte et al. (2023) take this work a step further by classifying hallucination mitigation strategies for foundation models in particular, in generative text and multi-modal settings respectively. Zhang et al. (2023b) similarly tackle hallucination detection for language foundation models, but limit their definition of hallucinations to conflicts between generations and inputs, contexts, and facts. The authors also discuss other possible problems in generated outputs, like ambiguity and bias. In the domain of image captioning, Li et al. (2023b) provide examples of hallucinations and present a new metric to evaluate object hallucinations. More recent surveys, like the ones from Bai et al. (2024) and Liu et al. (2024b), dive deeper into hallucination mitigation methods for multi-modal foundation models applied to image captioning tasks, but lack a general definition for hallucinations that encompasses decision-making applications. The extensive survey from Huang et al. (2023) primarily focuses on methods for evaluating the trustworthiness of language models from the fronts of factuality and faithfulness. Their taxonomy also briefly touches on hallucinations in object detection using LVLMs, but like other surveys, ignores broader decision-making deployments. There are also other works that provide examples of current use cases of L(V)LMs in autonomous vehicles (Yang et al., 2024c) and robotics (Zeng et al., 2023; Zhang et al., 2023a). Wang et al. (2023) perform a deep analysis of the trustworthiness of a variety of foundation models and Chen and Shu (2024) provide a taxonomy of hallucinations within LLMs, but both exclude applications to general decision problems. To the best of our knowledge, we are the first to propose a general definition of hallucinations that can be flexibly tuned to any particular deployment setting, including commonly found applications to QA or information retrieval, and more recent developments in planning or control. Furthermore, there is no existing work that summarizes state of the art methods for hallucination detection and mitigation approaches within decision-making and planning tasks. Additionally, to the best of our knowledge, ours is the first review to provide guidelines for choosing and designing hallucination intervention algorithms across different application areas.
@@
In the remainder of this work, we discuss the current uses of foundation models for decision-making tasks in Section 2, define and provide examples of hallucinations in Section 3, identify current detection methods in Section 4, present guidelines in Section 5, and explore research directions in Section 6.
2. Foundation Models Making Decisions
@@
Originally coined by Bommasani et al. (2022), the term foundation models refers to models that are “trained on broad data at scale such that they can be adapted to a wide range of downstream tasks.” This approach is in contrast to works that design and train models on a smaller subset of data for the purpose of being deployed to a specific task (Yang et al., 2024d). The key difference is that foundation models undergo a pre-training procedure on a large-scale dataset containing information from a variety of possible deployment fields, through which they are expected to learn more general features and correspondences that may be useful at test-time on a broader set of tasks (Zhou et al., 2023; Zhao et al., 2023). Examples of existing pre-trained foundation models span language (Devlin et al., 2019; Brown et al., 2020; Touvron et al., 2023a), vision (Caron et al., 2021; Oquab et al., 2024; Kirillov et al., 2023), and multi-modal (Radford et al., 2021; Achiam et al., 2023) inputs. In this section, we give a brief overview of existing use cases for foundation models in robotics and autonomous vehicles, and we provide a discussion of other decision-making systems in Appendix A.2. We also succinctly point out hallucinations found in these works and leave a lengthier discussion in Section 3.2. Readers should refer to works from Cui et al. (2024)Yang et al. (2024c)Zeng et al. (2023), and Zhang et al. (2023a) for a deeper review of application areas.
2.1. Autonomous Driving
@@
For the autonomous vehicle domain, researchers have formulated the use of language foundation models as a fine-tuning and prompt engineering problem (Wen et al., 2024b, a). An external sub-system is usually designed with (1) a perception module to process signals from raw sensors, (2) a memory bank of prior important experiences and its corresponding similarity function to find alike scenarios, and (3) a prompt generator to convert current sensor data and relevant memories into natural language that can be input to the foundation model. Currently, works either fine-tune LLMs with a few examples, or directly apply the model in a zero-shot manner, on a QA task with driving related questions. By framing the task in a QA form, researchers have been able to provide context to the L(V)LM to probe for high-level natural language decisions (Wen et al., 2024b, a), path planning (Mao et al., 2024; Sima et al., 2023), vehicle tracking and trajectory prediction (Wu et al., 2023b; Keysan et al., 2023), descriptions of the surroundings of the vehicle (Chen et al., 2024b; Xu et al., 2024), and low-level control (Liu et al., 2023). Figure 1 is an example of how a foundation model may be used in an autonomous driving setting, with possible hallucinations in deployment.
High-level Decisions
@@
Wen et al. (2024a) propose DiLu, a framework consisting of reasoning, reflection, and memory modules that support an LLM in producing high-level decisions for autonomous driving, and they test their method within a driving simulator environment. Specifically, the reasoning module views the current observation of the vehicle, queries the memory module for any similar situations that were encountered in the past, and converts the experience into a prompt, which is input to the LLM. The prompt is formatted such that it elicits chain-of-thought reasoning (Wei et al., 2022) from the LLM, which is shown to improve the accuracy of the model. The generated text output by the LLM is summarized by the reflection module, and is used to update the memory bank of experiences. A separate decision decoder model converts the summary into a discrete high-level decision (e.g., idle, turn right, accelerate, etc.) to take in the simulator. The same authors have also experimented with prompting the LVLM model, GPT-4V (Achiam et al., 2023), in a zero-shot manner to describe the surroundings of the vehicle, take high-level decisions, and explain why it believes it would be a good action to take (Wen et al., 2024b). They find that GPT-4V is capable of identifying safety-critical scenarios and suggests driving more conservatively in those situations. However, like other existing vision models (Suganuma and Yoneda, 2022), it has difficulty in detecting traffic light states. We discuss other hallucination examples in traffic scenarios in Section 3.2.
Figure 1. Example deployment of an LVLM foundation model in an autonomous driving setting. Hallucinations (pink) may arise at any point in the decision-making pipeline, including information retrieval, planning, perception, and control. In this example, the LVLM correctly queries the map for possible destinations of gas stations, but lists a location that is no longer open. Then, when navigating to one of the locations, the model predicts a path on the map that goes in the wrong direction of traffic flow. Finally, when applied to perception tasks for detecting possible pedestrians in front of the vehicle, the model hallucinates nearby people, causing an improper control action.
Path Planning
@@
Agent-Driver from Mao et al. (2024) utilizes a tool library of functions that communicate with neural modules that are responsible for object detection, trajectory prediction, occupancy estimation, and mapping. The LLM is asked to reason about what information would be helpful to plan a path of the ego vehicle, and calls on functions from the tool library to build up relevant context. Like Wen et al. (2024a), the authors use a memory bank of prior driving experiences to bolster the context provided to the LLM. With this context, the LLM predicts a suitable path for the ego vehicle to follow. If a collision is detected between the predicted trajectory and surrounding objects in the scene, the LLM undergoes self-reflection, like Reflexion (Shinn et al., 2023), another hallucination mitigation technique, to fine-tune its prediction. Through a short study, the authors test the frequency of invalid, hallucinated outputs from the model, and find that their self-reflection approach results in zero invalid generations at test-time. Sima et al. (2023) build up context before predicting a path for the ego vehicle by asking a VLM questions about its perception of surrounding vehicles, predicting their behaviors, planning high-level vehicle decisions, converting to lower level discrete actions, and finally, estimating a coordinate-level trajectory. Their method, DriveLM-Agent, predicts paths from images in an end-to-end manner using a multi-modal approach, whereas Agent-Driver requires sensor modules to process context separately.
Trajectory Prediction
@@
Works using foundation models for generating scene descriptions given multi-modal information frame the task as a QA problem. For example, Chen et al. (2024b) use a reinforcement learning (RL) agent pre-trained on the driving task in simulation to collect a dataset containing the vehicle’s state, environment observation (assumed to be ground truth from simulator,) low-level action, and the ego’s percentage of attention placed on different surrounding agents (if using an attention-based network). The authors ask ChatGPT (OpenAI, 2022) to act as a professional driving instructor who generates questions and answers given general driving rules and the vectorized description of the scene from the RL agent. Then, combining a pre-trained LLM with a vector embedder and former model, the architecture is trained end-to-end to answer questions from the scene. Examples of questions the architecture is posed at test-time include, “What objects are you observing,” “How are you going to drive in this situation and why,” and, “What are the best tourist spots in London” (the last of which is considered out of scope by the driving model). The authors acknowledge that the LLM may generate undesired hallucinated outputs during deployment, so they augment their training instruction dataset with out of scope questions that the model should learn to refuse to answer. DriveGPT4, proposed by Xu et al. (2024), is a multi-modal LLM using LLaMA 2 (Touvron et al., 2023b) which encodes video sequences with an encoder model, and projects question and video embeddings into a text form to be input to the LLM. A decoder model converts the tokenized output of the LLM into a low-level action with a corresponding high-level reason to take the action. Like Chen et al., the authors collect a driving QA dataset to align and fine-tune the architecture to driving tasks.
Control
@@
Liu et al. (2023) tackle low-level control within unsignalized intersections by formulating the problem as a multi-task decision-making process. Specifically, they train expert policies with RL to perform individual control tasks (e.g., turning left, going forward, and turning right), with which they collect an expert demonstration dataset. An LLM model, GPT-2 (Radford et al., 2019), is fine-tuned to predict next actions given past trajectories of state and action pairs from the expert dataset. The authors showcase that their method, MTD-GPT, is able to achieve higher success rates across each of the control tasks over the original RL policies alone, showcasing the promise of foundation models to leverage their general knowledge in a multi-task problem setting.
2.2. Robotics